The demand function for rose isa Q=a-bp and the supply function is Q=c+ep+ft, where a,b,c,e and for are positive constant and the is an average temperature in a month. Show how the equilibrium quantity and price vary with temperature?
1
Expert's answer
2015-06-09T11:21:04-0400
If demand function for rose is Qd = a­p-b and the supply function is Qs = c+ep+ft, then in equilibrium Qd = Qs, so: 
a­pb = c+ep+ft
p*(e + b) = a - c - ft
p = (a - c - ft)/(e + b)
So, equilibrium price increases, when the temperature falls and equilibrium quantity falls too.